# Estimation of Energy-dissipation Lower-bounds for Neuromorphic Learning-in-memory

### v3

@@ Results
Figure 6: Energy dissipation estimates when training using different LIM variants(A) Estimated instantaneous power dissipation $P$ and energy barrier height $E_{n}^{0}$ for training brain-scale AI using LIMB variants with a polynomial update rate $J_{n}\sim\frac{1}{n^{1+\gamma}}$ with $\{P1,E_{n}^{0}1\sim\gamma=0.5;P2,E_{n}^{0}2\sim\gamma=2;P3,E_{n}^{0}3\sim\gamma=5;P4,E_{n}^{0}4\sim\gamma=10\}$; (B) Total energy consumption of LIM under update rate schedules in (A): $E_{total}1=$$1.70\text{\times}{10}^{12}\text{\,}\mathrm{J}$; $E_{total}2=$$2.53\text{\times}{10}^{9}\text{\,}\mathrm{J}$; $E_{total}3=$$2.68\text{\times}{10}^{12}\text{\,}\mathrm{J}$; $E_{total}4=$$2.70\text{\times}{10}^{12}\text{\,}\mathrm{J}$; (C) Total LIM energy dissipation against update rate hyperparameter $\gamma$ in discrete finite simulation and infinite Riemann-zeta forms ; (D) Estimated energy dissipated compared against the reported energy dissipation for real-world AI training workloads [60]. From left to right, the corresponding AI models are T5, Meena, Gshard-600B, Switch Transformer, GPT-3, and LaMDA. The Brain-scale AI model estimates is for a precision $\delta=2^{-16}$. LIMA estimate is based on Eq. 22, and estimate of CEB limit is based on Eq. 19.
@@
Now we shift focus to explore the fundamental energy bound for training such workloads under our LIM theoretical framework. Fig. 6(A) plots the energy barrier-height $E_{n}^{0}1-E_{n}^{0}4$ and the estimated instantaneous power $\Delta E_{n}J_{n}$ with different update-rate schedules P1P4P1-P4. Each schedule corresponds to a different value of $\gamma$ in $r_{n}=1/n^{1+\gamma}$. Note that for this experiment we have used the numerical expression given by Eq. 18 for parameters #FLOPS\#FLOPS and $M$ that correspond to a brain-scale AI workload (extrapolated from Fig 5). For the polynomially decreasing update rate, the barrier height increases logarithmically. As a result, the instantaneous power dissipation for LIM systems decreases monotonically. Also note that at the end of the training $E^{0}_{\infty}$ is higher and satisfies the constraint 20.
@@ Discussion
In this paper we presented the energy-dissipation estimates for LIM that are model-agnostic and depends only on the training #FLOPS\#FLOPS, the number of model parameters $M$, the precision of memory retention $\delta$ and a hyperparameter $\gamma$ that models the decay in the update-rate schedule. In particular Eq. 18 describes how the memory energy barrier height is connected to two important parameters: (a) the parameter update rate; and (b) the learning rate, both of which determine two of the three performance walls, namely, the update-wall and the consolidation-wall. For instance, the update-wall is reflected in the profile of the update-rate $J_{n}$ for each of the parameters, and Eq. 18 shows how a specific update-profile $J_{n}$ can be achieved by modulating the barrier-profile in the learning-in-memory paradigm. Similarly, the learning rate $\epsilon_{n}$ determines the consolidation-wall. Several adaptive synaptic models have been proposed [29, 30] that show how a specific learning-rate profile can lead to optimal information transfer rate between short-term and long-term memories. In the LIM paradigm, the mechanism by which external energy is dissipated to drive the memory/synaptic updates determines different LIM variants. Like a Lyapunov system, the network or the algorithmic gradient is source of energy and determines the dynamics. From a thermodynamic computing point of view, the system (or memory) is initialized in a low-entropy state and is guided towards the final state through thermal fluctuations and by energy-barrier modulation. Thus, like Fig. 4(F), the energy, memory and compute functions are all integrated, a concept originally proposed by Feynmann and Bennett [67, 46]. Unlike a traditional computing that requires periodic resets and where entropy production leads to unavoidable energy-loss [41, 38], the LIM approach is a single-shot computation without any resets. However, in estimating energy-dissipation for LIM, we have ignored some practical constraints which we point out here. We have assumed that the energy-barrier modulation trajectory is pre-determined which then leads to the desired final state (result of the computation). In practice, it might not be possible to achieve this and would require dissipating additional energy in learning or adapting the modulation procedure. Secondly, we have assumed that the energy used to change the energy-barrier height can be recovered which in practice, could be challenging. We have also assumed that the reservoirs driving the memory elements do not dissipate energy which in practice is also difficult to achieve. Larger values of $\gamma$ or a faster update-rate decay implies that a higher $J_{max}$ is needed to satisfy the computational constraint for the training to terminate, i.e., the update-rate $J_{n}$ should also decay asymptotically to zero. In practice, the maximum update rate $J_{max}$ is determined by the maximum switching frequency ($f_{T}$) of a physical switching devices. For example, silicon-germanium heterojunction bipolar transistor [68] and an ultrafast optical switch [69] can exhibit maximum switch-rates close to $1\text{\,}\mathrm{THz}$, equivalently, $J_{max}\approx 10^{12}s^{-1}$.
@@
Irrespective of the energy dissipation mechanisms, all LIM variants are based on modulating the memory energy barrier according to Eq. 18 and Fig. 3. Energy barrier modulation supporting the LIM paradigm could be implemented in a variety of physical substrates using emerging memory devices. For instance, recently, we reported a dynamic memory device [35] that could also be used to modulate the memory retention profile and could be an attractive candidate to implement the LIM paradigm. However, note that to approach the fundamental energy limits of training/learning one would need to address all three performance walls. Compute-in-memory (CIM) alternatives where the computation and memory are vertically integrated in massively parallel, distributed architecture offer substantially greater computational bandwidth and energy efficiency in memristive neuromorphic cognitive computing [13] approaching the nominal energy efficiency of synaptic transmission in the human brain [70]. Resonant adiabatic switching techniques in charge-based CIM [71] further extend the energy efficiency by recycling the energy required to move charge by coupling the capacitive load to an inductive tank at resonance, providing a path towards efficiencies in cognitive computing superior to biology and, in principle, beyond the Landauer limit by overcoming the constraints of irreversible dissipative computing. It is an open question whether the learning-in-memory energy bounded by Eq. 8 could also be at least partially recovered through principles of adiabatic energy recycling.
@@
The code for simulation and data that support the findings of this article are openly available at [76].
@@ A.2 Calibration of LIM hyperparameters
Here we relate the model-specific hyperparameters $\beta$, $\lambda$, and $c$ in Eq. 16 to an LIM parameter that is model agnostic. For gradient-based learning the parameter $\epsilon_{n}>0$ needs to satisfy the following dynamic constraint:
@@
While many time-schedules for $\epsilon_{n}$ satisfying the constaints 25 are admissible, an $\epsilon_{n}=\mathcal{O}\left({1/\sqrt{n}}\right)$ schedule has been proposed [29, 54, 36] for achieving optimal memory consolidation in benchmark random pattern experiments. From Eq. 12, substituting $\epsilon_{n}$ for the optimal schedule,
$\begin{split}\Delta E_{n}&=\beta kT\sigma_{g}^{2}\sum_{k=1}^{n}\prod_{m=k+1}^{n}(1-\lambda\epsilon_{m})\epsilon_{k}^{2}\\ &=\beta kT\sigma_{g}^{2}\sum_{k=1}^{n}\exp\left(\sum_{m=k+1}^{n}\log(1-\lambda\epsilon_{m})\right)\epsilon_{k}^{2}\end{split}$ (26)
@@
[…]
@@
$\begin{split}\Delta E_{n}&\geq\beta kT\sigma_{g}^{2}\frac{c^{2}}{n}\sum_{p=0}^{n-1}\exp\left(-\lambda c\left(\sqrt{n}-\sqrt{n}\sqrt{1-\frac{p}{n}}\right)\right)\\ &\geq\beta kT\sigma_{g}^{2}\frac{c^{2}}{n}\sum_{k=1}^{n}\exp\left(-2\lambda c\frac{p}{\sqrt{n}}\right)\\ &\geq\beta kT\sigma_{g}^{2}\frac{c^{2}}{n}\frac{1}{1-\exp{\left(-\frac{2\lambda c}{\sqrt{n}}\right)}}\quad(\text{Geometric series}),\end{split}$ (28)
@@
$\begin{split}\Delta E_{n}&\geq\beta kT\sigma_{g}^{2}\frac{c^{2}}{n}\frac{\sqrt{n}}{2\lambda c}=C/\sqrt{n},\end{split}$ (29)
@@
Again, numerous choices for scheduling $J_{n}$ to satisfy the constraint 30 are possible. Since the update rate $J_{n}$ proceeds at least as fast as the learning rate $\epsilon_{n}$, to achieve the asymptotic bound, we make $J_{n}:=\epsilon_{n}$. Note that for this condition the total number of updates becomes unbounded or
@@ A.3 LIM energy estimates for large AI models
For a learning rate schedule $\epsilon_{n}=1/\sqrt{n}$, the update-rate schedule $r_{n}\sim\frac{J_{n}}{J_{max}}=1/n^{1+\gamma}$ with $\gamma>0$ ensures that Eq. 30and 20 are satisfied. In LIMA, the energy to drive the memory updates are derived from the algorithmic gradient. Therefore, according to Eq. 33 and the asymptotic calibration, $\Delta E_{n}^{diss}=\epsilon_{n}\left(kT/\delta\right)$, Eq. 22 leads to
@@ A.4 LIM analysis for exponential decay update-rate
$\begin{split}E_{Total}&\approx M\left(\frac{kT}{\delta}\sum_{n=1}^{\infty}{R_{n}}\epsilon_{n}\Delta t+E_{\infty}^{0}\right)\\ &=\#FLOPs\left(\frac{kT}{\delta}\right)\left[\frac{\sum_{n=1}^{\infty}\frac{1}{n}e^{-\gamma n}}{\sum_{n=1}^{\infty}e^{-\gamma n}}\right]+ME_{\infty}^{0}\\ &=\#FLOPs\left(\frac{kT}{\delta}\right)\left[(1-e^{\gamma})\log{\left(\frac{e^{\gamma n}}{e^{\gamma}-1}\right)}\right]\\ &+MkT\log\left[\frac{1}{\delta}\right]\end{split}$ (36)

### v4

@@ Results
Figure 6: Energy dissipation estimates when training under LIM paradigm(A) Estimated instantaneous power dissipation $P$ and energy barrier height $E_{n}^{0}$ for training brain-scale AI using LIM lower-bounds with a polynomial update rate $J_{n}\sim\frac{1}{n^{1+\gamma}}$ with $\{P1,E_{n}^{0}1\sim\gamma=0.5;P2,E_{n}^{0}2\sim\gamma=2;P3,E_{n}^{0}3\sim\gamma=5;P4,E_{n}^{0}4\sim\gamma=10\}$; (B) Total energy consumption of LIM under update rate schedules in (A): $E_{total}1=$$1.70\text{\times}{10}^{12}\text{\,}\mathrm{J}$; $E_{total}2=$$2.53\text{\times}{10}^{9}\text{\,}\mathrm{J}$; $E_{total}3=$$2.68\text{\times}{10}^{12}\text{\,}\mathrm{J}$; $E_{total}4=$$2.70\text{\times}{10}^{12}\text{\,}\mathrm{J}$; (C) Total LIM energy dissipation against update rate hyperparameter $\gamma$ in discrete finite simulation and infinite Riemann-zeta forms ; (D) Estimated energy dissipated compared against the reported energy dissipation for real-world AI training workloads [58]. From left to right, the corresponding AI models are T5, Meena, Gshard-600B, Switch Transformer, GPT-3, and LaMDA. The Brain-scale AI model estimates is for a precision $\delta=2^{-16}$. LIMA estimate is based on Eq. 14, and estimate of CEB limit is based on Eq. 12.
@@
Now we shift focus to explore the fundamental energy bound for training such workloads under our LIM theoretical framework. Fig. 6(A) plots the energy barrier-height $E_{n}^{0}1-E_{n}^{0}4$ and the estimated instantaneous power $\Delta E_{n}J_{n}$ with different update-rate schedules P1P4P1-P4. Each schedule corresponds to a different value of $\gamma$ in $r_{n}=1/n^{1+\gamma}$. Note that for this experiment we have used the numerical expression given by Eq. .2 for parameters #FLOPS\#FLOPS and $M$ that correspond to a brain-scale AI workload (extrapolated from Fig 5). For the polynomially decreasing update rate, the barrier height increases logarithmically. As a result, the instantaneous power dissipation for LIM systems decreases monotonically. Also note that at the end of the training $E^{0}_{\infty}$ is higher and satisfies the constraint 12.
@@ Discussion
In this paper we presented the energy-dissipation estimates for LIM that are model-agnostic and depends only on the training #FLOPS\#FLOPS, the number of model parameters $M$, the precision of memory retention $\delta$ and a hyperparameter $\gamma$ that models the decay in the update-rate schedule. In particular Eq. .2 describes how the memory energy barrier height is connected to two important parameters: (a) the parameter update rate; and (b) the learning rate, both of which determine two of the three performance walls, namely, the update-wall and the consolidation-wall. For instance, the update-wall is reflected in the profile of the update-rate $J_{n}$ for each of the parameters, and Eq. .2 shows how a specific update-profile $J_{n}$ can be achieved by modulating the barrier-profile in the learning-in-memory paradigm. Similarly, the learning rate $\epsilon_{n}$ determines the consolidation-wall. Several adaptive synaptic models have been proposed [23, 43] that show how a specific learning-rate profile can lead to optimal information transfer rate between short-term and long-term memories. In the LIM paradigm, the mechanism by which external energy is dissipated to drive the memory/synaptic updates is determined by the gradients provided by the learning algorithm. Like a Lyapunov system, the network or the algorithmic gradient is source of energy and determines the dynamics. From a thermodynamic computing point of view, the system (or memory) is initialized in a low-entropy state and is guided towards the final state through thermal fluctuations and by energy-barrier modulation. Thus, like Fig. 4(F), the energy, memory and compute functions are all integrated, a concept originally proposed by Feynmann and Bennett [20, 4]. Unlike a traditional computing that requires periodic resets and where entropy production leads to unavoidable energy-loss [74, 75], the LIM approach is a single-shot computation without any resets. However, in estimating energy-dissipation for LIM, we have ignored some practical constraints which we point out here. We have assumed that the energy-barrier modulation trajectory is pre-determined which then leads to the desired final state (result of the computation). In practice, it might not be possible to achieve this and would require dissipating additional energy in learning or adapting the modulation procedure. (Secondly, we have assumed that the energy used to change the energy-barrier height is unrecoverable, resulting in the static term in Eq. 12. In practice, the energy dissipated due to barrier modulation could be recovered but is challenging and would incur additional energy cost. We have also assumed that the reservoirs driving the memory elements do not dissipate energy which in practice is also difficult to achieve. Larger values of $\gamma$ or a faster update-rate decay implies that a higher $J_{max}$ is needed to satisfy the computational constraint for the training to terminate, i.e., the update-rate $J_{n}$ should also decay asymptotically to zero. In practice, the maximum update rate $J_{max}$ is determined by the maximum switching frequency ($f_{T}$) of a physical switching devices. For example, silicon-germanium heterojunction bipolar transistor [13] and an ultrafast optical switch [36] can exhibit maximum switch-rates close to $1\text{\,}\mathrm{THz}$, equivalently, $J_{max}\approx 10^{12}s^{-1}$.
(From the view point of learning, we assumed in our derivations that the learning gradients are temporally uncorrelated with a constant variance. This is an idealization for obtaining an analytically tractable lower-bound, as described in Eq.10. In practice, many commonly used learning mechanisms are known to induce temporal gradient correlations and violate the statistical independence assumptions. Examples of such learning algorithms include momentum and Adam optimizers [60, 42], learning-rate schedules and restart schemes [50, 69], stateful normalization layers [37], data-sampling effects [48] and curvature-induced nonstationarity [76]. These correlations could be exploited to reduce the energy dissipation bounds even further, however the analysis would require Monte-Carlo based estimation.
@@
Irrespective of the energy dissipation mechanisms, LIM energy dissipation lower bounds are based on modulating the memory energy barrier according to Eq. .2 and Fig. 3. Energy barrier modulation supporting the LIM paradigm could be implemented in a variety of physical substrates using emerging memory devices. For instance, recently, we reported a dynamic memory device [54] that could also be used to modulate the memory retention profile and could be an attractive candidate to implement the LIM paradigm. However, note that to approach the fundamental energy limits of training/learning one would need to address all three performance walls. Compute-in-memory (CIM) alternatives where the computation and memory are vertically integrated in massively parallel, distributed architecture offer substantially greater computational bandwidth and energy efficiency in memristive neuromorphic cognitive computing [72] approaching the nominal energy efficiency of synaptic transmission in the human brain [10]. Resonant adiabatic switching techniques in charge-based CIM [39] further extend the energy efficiency by recycling the energy required to move charge by coupling the capacitive load to an inductive tank at resonance, providing a path towards efficiencies in cognitive computing superior to biology and, in principle, beyond the Landauer limit by overcoming the constraints of irreversible dissipative computing. It is an open question whether the learning-in-memory energy bounded by Eq. 6 could also be at least partially recovered through principles of adiabatic energy recycling.
@@
The code for simulation and data that support the findings of this article are publicly available at [16].
@@ A.2 Calibration of LIM hyperparameters
Here we relate the model-specific hyperparameters $\beta$, $\lambda$, and $c$ in Eq. .2 to an LIM parameter that is model agnostic. For gradient-based learning the parameter $\epsilon_{n}>0$ needs to satisfy the following dynamic constraint:
@@
While many time-schedules for $\epsilon_{n}$ satisfying the constaints 17 are admissible, an $\epsilon_{n}=\mathcal{O}\left({1/\sqrt{n}}\right)$ schedule has been proposed [23, 3, 61] for achieving optimal memory consolidation in benchmark random pattern experiments. From Eq. 10, substituting $\epsilon_{n}$ for the optimal schedule,
$split\Delta E_{n}&=\beta kT\sigma_{g}^{2}\sum_{k=1}^{n}\prod_{m=k+1}^{n}(1-\lambda\epsilon_{m})\epsilon_{k}^{2}\\ &=\beta kT\sigma_{g}^{2}\sum_{k=1}^{n}\exp\left(\sum_{m=k+1}^{n}\log(1-\lambda\epsilon_{m})\right)\epsilon_{k}^{2}$ (18)
@@
$split\Delta E_{n}&=\beta kT\sigma_{g}^{2}\sum_{k=1}^{n}\exp\left(-\lambda c(\sqrt{n}-\sqrt{k})\right)\left(\frac{c^{2}}{k}\right)\\ &\geq\beta kT\sigma_{g}^{2}\frac{c^{2}}{n}\sum_{k=1}^{n}\exp\left(-\lambda c(\sqrt{n}-\sqrt{k})\right)\\$ (19)
@@
$split\Delta E_{n}&\geq\beta kT\sigma_{g}^{2}\frac{c^{2}}{n}\sum_{p=0}^{n-1}\exp\left(-\lambda c\left(\sqrt{n}-\sqrt{n}\sqrt{1-\frac{p}{n}}\right)\right)\\ &\geq\beta kT\sigma_{g}^{2}\frac{c^{2}}{n}\sum_{k=1}^{n}\exp\left(-2\lambda c\frac{p}{\sqrt{n}}\right)\\ &\geq\beta kT\sigma_{g}^{2}\frac{c^{2}}{n}\frac{1}{1-\exp{\left(-\frac{2\lambda c}{\sqrt{n}}\right)}}\quad(\text{Geometric series}),$ (20)
@@
$split\Delta E_{n}&\geq\beta kT\sigma_{g}^{2}\frac{c^{2}}{n}\frac{\sqrt{n}}{2\lambda c}=C/\sqrt{n},$ (21)
@@
Again, numerous choices for scheduling $J_{n}$ to satisfy the constraint 22 are possible. Since the update rate $J_{n}$ proceeds at least as fast as the learning rate $\epsilon_{n}$, to achieve the asymptotic bound, we make $J_{n}:=\epsilon_{n}$. Note that for this condition the total number of updates becomes unbounded or
@@ A.3 LIM energy estimates for large AI models
For a learning rate schedule $\epsilon_{n}=1/\sqrt{n}$, the update-rate schedule $r_{n}\sim\frac{J_{n}}{J_{max}}=1/n^{1+\gamma}$ with $\gamma>0$ ensures that Eq. 22and 12 are satisfied. In LIMA, the energy to drive the memory updates are derived from the algorithmic gradient. Therefore, according to Eq. 25 and the asymptotic calibration, $\Delta E_{n}^{diss}=\epsilon_{n}\left(kT/\delta\right)$, Eq. 14 leads to
@@ A.4 LIM analysis for exponential decay update-rate
$split E_{Total}&\approx M\left(\frac{kT}{\delta}\sum_{n=1}^{\infty}{R_{n}}\epsilon_{n}\Delta t+E_{\infty}^{0}\right)\\ &=\#FLOPs\left(\frac{kT}{\delta}\right)\left[\frac{\sum_{n=1}^{\infty}\frac{1}{n}e^{-\gamma n}}{\sum_{n=1}^{\infty}e^{-\gamma n}}\right]+ME_{\infty}^{0}\\ &=\#FLOPs\left(\frac{kT}{\delta}\right)\left[(1-e^{\gamma})\log{\left(\frac{e^{\gamma n}}{e^{\gamma}-1}\right)}\right]\\ &+MkT\log\left[\frac{1}{\delta}\right]$ (28)